OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = −three fourthsx + 8 y = −three fourthsx + 3

OpenStudy (aroub):

4x-3y=-12 -3y=-4x-12 y=4/3x+4 Slope= 4/3 and remember parallel=same slope y=mx+b , m=4/3 , (-3,4) 4=4/3(-3)+b 4=-4+b 8=b => y=4/3x+8

OpenStudy (anonymous):

THNK SO MUCH:)

OpenStudy (aroub):

YOU'RE WELCOME :)

OpenStudy (anonymous):

CAN U HELP ME WITH ANOTHR ONE PLZ??

OpenStudy (aroub):

YES!

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = one thirdx − 3 y = one thirdx + 17 y = −3x − 3 y = −3x + 17

OpenStudy (aroub):

m=1/3 , (6,-1) y=mx+b -1=1/3(6)+b -1=2+b -3=b => y=1/3x-3

OpenStudy (anonymous):

thank you:)

OpenStudy (aroub):

No problem :)